7x + 4y = -4
5x + 8x = 28
how do you solve the system of equations by elimination

Answer by persian52(161)

7x+4y=-4_5x+8x=28
►Since 5x and 8x are like terms, add 8x to 5x to get 13x.
7x+4y=-4_13x=28
►Divide each term in the equation by 13.
7x+4y=-4_(13x)/(13)=(28)/(13)
►Simplify the left-hand side of the equation by canceling the common terms.
7x+4y=-4_x=(28)/(13)
►Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is (28)/(13).
7((28)/(13))+4y=-4_x=(28)/(13)
►Multiply 7 by each term inside the parentheses.
(196)/(13)+4y=-4_x=(28)/(13)
►Since (196)/(13) does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting (196)/(13) from both sides.
4y=-(196)/(13)-4_x=(28)/(13)
►Simplify the right-hand side of the equation.
4y=-(248)/(13)_x=(28)/(13)
►Divide each term in the equation by 4.
(4y)/(4)=-(248)/(13)*(1)/(4)_x=(28)/(13)
►Simplify the left-hand side of the equation by canceling the common terms.
y=-(248)/(13)*(1)/(4)_x=(28)/(13)
►Simplify the right-hand side of the equation by simplifying each term.
y=-(62)/(13)_x=(28)/(13)
►Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is -(62)/(13).
=► y=-(62)/(13)
=► x=(28)/(13)